# VASA-1: Lifelike Audio-Driven Talking Faces Generated in Real Time

### v1

@@ Abstract
We introduce VASA, a framework for generating lifelike talking faces with appealing visual affective skills (VAS) given a single static image and a speech audio clip. Our premiere model, VASA-1, is capable of not only producing lip movements that are exquisitely synchronized with the audio, but also capturing a large spectrum of facial nuances and natural head motions that contribute to the perception of authenticity and liveliness. The core innovations include a diffusion-based holistic facial dynamics and head movement generation model that works in a face latent space, and the development of such an expressive and disentangled face latent space using videos. Through extensive experiments including evaluation on a set of new metrics, we show that our method significantly outperforms previous methods along various dimensions comprehensively. Our method delivers high video quality with realistic facial and head dynamics and also supports the online generation of 512$\times$512 videos at up to 40 FPS with negligible starting latency. It paves the way for real-time engagements with lifelike avatars that emulate human conversational behaviors. Project webpage: https://www.microsoft.com/en-us/research/project/vasa-1/
@@ 1 Introduction
To achieve our goal, another challenge lies in constructing the latent space for the aforementioned holistic facial dynamics and gathering the data for the diffusion model training. Beyond facial and head movements, a human face image contains other factors such as identity and appearance. In this work, we seek to build a proper latent space for human face using a large volume of face videos . Our aim is for the face latent space to possess both a total state of disentanglement between facial dynamics and other factors, as well as a high degree of expressiveness to model rich facial appearance details and dynamic nuances. We base our method on the 3D-aided representation [61, 17] and equip it with a collection of carefully-designed loss functions. Trained on face videos in an self-supervised or weakly-supervised manner, our encoder can produce well-disentangled factors including 3D appearance, identity, head pose and holistic facial dynamics, and the decoder can generate high quality faces following the given latent codes.
VASA-1 has collectively advanced the realism of lip-audio synchronization, facial dynamics, and head movement to new heights. Coupled with high image generation quality and efficient running speed, we achieved real-time talking faces that are realistic and lifelike. Through detailed evaluations, we show that our method significantly outperforms existing methods. We believe VASA-1 brings us closer to a future where digital AI avatars can engage with us in ways that are as natural and intuitive as interactions with real humans, demonstrating appealing visual affective skills for more dynamic and empathetic information exchange.
@@ 2 Related Work
The representation of facial images through disentangled variables has been extensively studied by previous works. Some methods utilize sparse keypoints [42, 69] or 3D face models [40, 20, 71] to explicitly characterize facial dynamics and other properties, but these can suffer from issues such as inaccurate reconstructions or limited expressive capabilities. There are also many works dedicated to learning disentangled representations within a latent space. A common approach involves separating faces into identity and non-identity components, then recombining them across different frames, either in a 2D [10, 74, 32, 67, 35, 57] or 3D context [61, 17]. The main challenge faced by these methods is the effective disentanglement of various factors while still achieving expressive representations of all static and dynamic facial attributes, which is addressed in this work.
@@
Recent advances in generative models [9, 25, 46, 45] have led to significant progress in video generation. Earlier video generation approaches [56, 53, 44] employed the adversarial learning [22] framework, while more recent methods [66, 6, 21, 30, 3, 8] have leveraged diffusion or auto-regressive models to capture diverse video distributions. Recently, several works concurrent to us [52, 62] have adapted video diffusion techniques to audio-driven talking face generation, achieving promising results despite the slow training and inference speeds. In contrast, our method can deliver both efficiency and high-quality results in the generation of talking face videos.
@@ 3 Method
Task definition.
As illustrated in Fig. 1, The input to our method consists of a single face image $\mathbf{I}$ of an arbitrary identity and a speech audio clip $\mathbf{a}$ from, again, an arbitrary person. The goal is to generate a synthesized video of the input face image speaking with the given audio in a realistic and coherent manner. A successfully generated video should exhibit high fidelity in several key aspects: the clarity and authenticity of the image frames, precise synchronization between the audio and lip movements, expressive and emotive facial dynamics, and naturalistic head poses.
Our generation process can also accept a set of optional control signals to guide the generation, which include the main eye gaze direction $\mathbf{g}$, head-to-camera distance $d$, and emotion offset $\mathbf{e}$. More details will be provided in the later sections.
@@
Instead of generating video frames directly, we generate holistic facial dynamics and head motion in the latent space conditioned on audio and other signals. Given these motion latent codes, our method then produces video frames by a face decoder, which also takes the appearance and identity features extracted using a face encoder from the input image as input.
To achieve this, we start by constructing a face latent space and training the face encoder and decoder. An expressive and disentangled face latent learning framework is crafted and trained on real-life face videos. Then we train a simple yet powerful Diffusion Transformer to model the motion distribution and generate the motion latent codes in the test time given audio and other conditions.
@@ 3.1 Expressive and Disentangled Face Latent Space Construction
To learn the disentangled latent space, the core idea is to construct image reconstruction loss by swapping latent variables between different images in videos. Our basic loss functions are adapted from [17]. However, we identified the poor disentanglement between facial dynamics and head pose using the original losses. The disentanglement between identity and motions is also imperfect. Therefore, we introduce several additional losses crucial to achieve our goal. For instance, inspired by [35], we add a pairwise head pose and facial dynamics transfer loss to improve their disentanglement. Let $\mathbf{I}_{i}$ and $\mathbf{I}_{j}$ be two frames randomly sampled from the same video of a subject. We extract their latent variables using the encoders, and transfer $\mathbf{I}_{i}$’s head pose onto $\mathbf{I}_{j}$ as $\hat{\mathbf{I}}_{j,\mathbf{z}^{pose}_{i}}=\mathcal{D}(\mathbf{V}^{app}_{j},\mathbf{z}^{id}_{j},\mathbf{z}^{pose}_{i},\mathbf{z}^{dyn}_{j})$ and $\mathbf{I}_{j}$’s facial motion onto $\mathbf{I}_{i}$ as $\hat{\mathbf{I}}_{i,\mathbf{z}^{dyn}_{j}}=\mathcal{D}(\mathbf{V}^{app}_{i},\mathbf{z}^{id}_{i},\mathbf{z}^{pose}_{i},\mathbf{z}^{dyn}_{j})$. The discrepancy loss between $\hat{\mathbf{I}}_{j,\mathbf{z}^{pose}_{i}}$ and $\hat{\mathbf{I}}_{i,\mathbf{z}^{dyn}_{j}}$ is subsequently minimized. To reinforce the disentanglement between identity and motions, we add a face identity similarity loss for the cross-identity pose and facial motion transfer results. Let $\mathbf{I}_{s}$ and $\mathbf{I}_{d}$ be the video frames of two different subjects, we can transfer the motions of $\mathbf{I}_{d}$ onto $\mathbf{I}_{s}$ and obtain $\hat{\mathbf{I}}_{s,\mathbf{z}^{pose}_{d},\mathbf{z}^{dyn}_{d}}=\mathcal{D}(\mathbf{V}^{app}_{s},\mathbf{z}^{id}_{s},\mathbf{z}^{pose}_{d},\mathbf{z}^{dyn}_{d})$. Then, a cosine similarity loss between the deep face identity features [15] extracted from $\mathbf{I}_{s}$ and $\hat{\mathbf{I}}_{s,\mathbf{z}^{pose}_{d},\mathbf{z}^{dyn}_{d}}$ is applied.
@@ 3.2 Holistic Facial Dynamics Generation with Diffusion Transformer
In order to achieve a seamless transition between adjacent windows, we incorporate the last $K$ frames of the audio feature and generated motions from the previous window as the condition of the current one. To summarize, our input condition can be denoted as $\mathbf{C}=[\mathbf{X}^{pre},\mathbf{A}^{pre};\mathbf{A},\mathbf{g},d,\mathbf{e}]$
@@ 4 Experiments
For face latent space learning, we use the public VoxCeleb2 dataset from [13] which contains talking face videos from about 6K subjects. We reprocess the dataset and discard the clips with multiple individuals and those of low quality using the method of [48]. For motion latent generation, we use an 8-layer transformer encoder with an embedding dim $512$ and head number $8$ as our diffusion network. The model is trained on VoxCeleb2 [13] and another high-resolution talk video dataset collected by us, which contains about 3.5K subjects. In our default setup, the model uses a forward-facing main gaze condition, an average head distance of all training videos, and an empty emotion offset condition. The CFG parameters are set to $\lambda_{\mathbf{A}}=0.5$ and $\lambda_{\mathbf{g}}=1.0$, and $50$ sampling steps are used.
@@
Figure 4: Disentanglement between identity and motion. In these examples, the same generated head and facial motion sequences are applied onto three different face images.
Figure 5: Disentanglement between head pose and facial dynamics. From top to bottom: the raw generated sequence, applying generated poses with fixed initial facial dynamics, and applying generated facial dynamics with fixed initial head pose and pre-defined spinning poses, respectively.
4.1 Qualitative Evaluation
Visual results.
Figure 1 presents some representative audio-driven talking face generation results of our method. Visually inspected, our method can generate high-quality video frames with vivid facial emotions. Moreover, it can generate human-like conversational behaviors, including sporadic shifts in eye gaze during speech and contemplation, as well as the natural and variable rhythm of eye blinking, among other nuances. We highly recommend that readers view our video results online to fully perceive the capabilities and output quality of our method.
Generation controllability.
Figure 3 shows our generated results under different control signals including main eye gaze, head distance, and emotion offset. It is evident that our generation model can well interpret these signals and produce talking face results that closely adhere to these specified parameters.
Disentanglement of face latents.
Figure 4 shows that when applying the same motion latent sequences onto different subjects, our method effectively maintains both the distinct facial movements and the unique facial identities. This indicates the efficacy of our method in disentangling identity and motion. Figure 5 further illustrates the effective disentanglement between head pose and facial dynamics. By holding one aspect constant and changing the other, the resulting images faithfully reflect the intended head and facial motions without interference.
Out-of-distribution generation.
Our method exhibits the capability to handle photo and audio inputs that fall outside the training distribution. For instance, as shown in Figure 6, it can handle artistic photos, singing audio clips (top two rows), and non-English speech (the last row). Notably, these data variants were not present in the training dataset.
Figure 6: Generation results with out-of-distribution images (non-photorealistic) and audios (singing audios for the first two rows and non-English speech for the last row). Our method can still generate high quality videos well-aligned with the audios, although it was not trained on such data variations.
@@ 4.2 Quantitative Evaluation
  • Pose variation intensity. We further define a pose variation intensity score $\Delta P$ which is the average of the pose angle differences between adjacent frames. Averaged over all the frames of all generated videos, $\Delta P$ provides an indication of the overall head motion intensity generated by a method.
  • Video quality. Following previous video generation works [66, 44], we use the Fréchet Video Distance (FVD) [54] to evaluate the generated video quality. We compute the FVD metric using sequences of 25 consecutive frames.
@@
We compare our method with there existing audio-driven talking face generation methods: MakeItTalk [75], Audio2Head [59, 72], and SadTalker [72]. MakeItTalk [75] employs an LSTM to convert audio into dynamic facial landmarks, then use the landmarks to animate a source image into a video sequence through either image warping or neural network based image translation. Audio2Head [59] uses a motion-aware recurrent network to translate audio into head poses, which, along with the original audio, are used to generate dense motion fields. The motion fields are further applied to the image features and the final output is then generated by a neural network. SadTalker [72] employs a VAE network to generate pose offsets from audio and a regression network to predict lip-only coefficients from audio features. A random variable is used for eye blink generation. This method can generate varied poses and eye blinks from identical audio inputs but only regress a deterministic pattern for other motions such as eyebrow, gaze and facial expressions.
@@
For each audio input, we generate a single video for deterministic approaches, i.e., MakeItTalk and Audio2Head. For SadTalker [72] and our method, we sample three videos for each audio and average the computed metrics. Since different pose representations are used by these methods, we re-extract the head poses from the generated frames to compute the pose-related metrics (i.e., CAPP and $\Delta P$). For the FVD metric, we use 2K 25-frame video clips of both the real videos and generated ones. For reference purpose, we also report the evaluated metrics of real videos.
Table 1 and Table 2 present the results on the VoxCeleb2 and OneMin-32 benchmarks. Note that we did not evaluate the FVD on VoxCeleb2 as its video quality is varied and often low. On both benchmarks, our method achieves the best results among all methods on all evaluated metrics. In terms of audio-lip synchronization scores ($S_{C}$ and $S_{D}$), our method outperforms all others by a wide margin. Note that our method yields better scores than real videos, which is due to effect of the audio CFG (see Sec. 4.3). Our generated poses are better aligned with the audios especially on the OneMin-32 benchmark, as reflected by the CAPP scores. The head movements also exhibit the highest intensity according to $\Delta P$, although there’s still a gap to the intensity of real videos. Our FVD score is significantly lower than others, demonstrating the much higher video quality and realism of our results.
Table 1: Comparison with previous methods on the VoxCeleb2 benchmark.
[…]
MakeItTalk [75] 4.176 15.513 -0.051 0.210
Audio2Head [59] 6.172 8.470 0.246 0.260
SadTalker [72] 5.843 8.813 0.441 0.275
Ours $\mathbb{8.841}$ $\mathbb{6.312}$ $\mathbb{0.468}$ $\mathbb{0.304}$
Real video 7.640 7.189 0.588 0.505
Table 2: Comparison with previous methods on the OneMin-32 benchmark.
[…]
MakeItTalk [75] -0.123 14.340 0.002 0.190 304.833
Audio2Head [59] 5.992 8.211 0.205 0.239 209.772
SadTalker [72] 5.501 8.850 0.383 0.252 214.507
Ours $\mathbb{7.957}$ $\mathbb{6.635}$ $\mathbb{0.465}$ $\mathbb{0.316}$ $\mathbb{105.884}$
Real video 7.192 7.254 0.559 0.405 29.245

### v2

@@ Abstract
We introduce VASA, a framework for generating lifelike talking faces with appealing visual affective skills (VAS) given a single static image and a speech audio clip. Our premiere model, VASA-1, is capable of not only generating lip movements that are exquisitely synchronized with the audio, but also producing a large spectrum of facial nuances and natural head motions that contribute to the perception of authenticity and liveliness. The core innovations include a diffusion-based holistic facial dynamics and head movement generation model that works in a face latent space, and the development of such an expressive and disentangled face latent space using videos. Through extensive experiments including evaluation on a set of new metrics, we show that our method significantly outperforms previous methods along various dimensions comprehensively. Our method delivers high video quality with realistic facial and head dynamics and also supports the online generation of 512$\times$512 videos at up to 40 FPS with negligible starting latency. It paves the way for real-time engagements with lifelike avatars that emulate human conversational behaviors. Project webpage: https://www.microsoft.com/en-us/research/project/vasa-1/
@@ 1 Introduction
To achieve our goal, another challenge lies in constructing the latent space for the aforementioned holistic facial dynamics and gathering the data for the diffusion model training. Beyond facial and head movements, a human face image contains other factors such as identity and appearance. In this work, we seek to build a proper latent space for human face using a large volume of face videos. Our aim is for the face latent space to possess both a total state of disentanglement between facial dynamics and other factors, as well as a high degree of expressiveness to model rich facial appearance details and dynamic nuances. We base our method on the 3D-aided representation [64, 19] which was proven to be expressive, and equip it with a collection of newly-designed loss functions critical to effective disentanglement. Without the new designs we can never reach a high quality of talking face generation, especially the liveliness with nuanced emotions. Trained on face videos in an self-supervised or weakly-supervised manner, our encoder can produce well-disentangled factors including 3D appearance, identity, head pose and holistic facial dynamics, and the decoder can generate high quality faces following the given latent codes.
VASA-1 has collectively advanced the realism of lip-audio synchronization, facial dynamics, and head movement to new heights. Coupled with high image generation quality and efficient running speed, we achieved real-time talking faces that are realistic and lifelike. Through detailed evaluations, we show that our method significantly outperforms existing methods on a set of metrics, including a novel data-driven metric called Contrastive Audio and Pose Pretraining (CAPP) for measuring the audio-pose alignment and a pose variation intensity score that is related to the vividness of head motion. We believe VASA-1 brings us closer to a future where digital AI avatars can engage with us in ways that are as natural and intuitive as interactions with real humans, demonstrating appealing visual affective skills for more dynamic and empathetic information exchange.
@@ 2 Related Work
The representation of facial images through disentangled variables has been extensively studied by previous works. Some methods utilize sparse keypoints [44, 72] or 3D face models [42, 22, 73] to explicitly characterize facial dynamics and other properties, but these can suffer from issues such as inaccurate reconstructions or limited expressive capabilities. There are also many works dedicated to learning disentangled representations within a latent space. A common approach involves separating faces into identity and non-identity components, then recombining them across different frames, either in a 2D [11, 76, 34, 70, 37, 60, 54] or 3D context [64, 19, 18]. The main challenge faced by these methods is the effective disentanglement of various factors while still achieving expressive representations of all static and dynamic facial attributes, which is addressed in this work.
@@
Recent advances in generative models [10, 27, 48, 47] have led to significant progress in video generation. Earlier video generation approaches [59, 56, 46] employed the adversarial learning [24] framework, while more recent methods [69, 7, 23, 32, 4, 9] have leveraged diffusion or auto-regressive models to capture diverse video distributions. Recently, several works concurrent to us [55, 65] have adapted video diffusion techniques to audio-driven talking face generation, achieving promising results despite the slow training and inference speeds. In contrast, our method is able to not only generating high-quality results but also achieve real-time efficiency – a metric crucial to efficiency-demanding applications such as live communication.
@@ 3 Method
As illustrated in Fig. 1, our method takes a single face image, optional control signals, and a speech audio clip to produce a realistic talking face video. Instead of generating video frames directly, we generate holistic facial dynamics and head motion in the latent space conditioned on audio and other signals. To achieve this, we start by constructing a face latent space and training the face encoder and decoder. An expressive and disentangled face latent learning framework is crafted and trained on real-life face videos. Then we train a simple yet powerful Diffusion Transformer to model the motion distribution and generate the motion latent codes in the test time given audio and other conditions.
@@ 3.1 Expressive and Disentangled Face Latent Space Construction
To learn the disentangled latent space, the core idea is to construct image reconstruction loss by swapping latent variables between different images in videos. Our basic loss functions are adapted from [19]. However, we identified the poor disentanglement between facial dynamics and head pose using the original losses. The disentanglement between identity and motions is also imperfect. Therefore, we introduce several additional losses crucial to achieve our goal. Inspired by [37], we add a pairwise head pose and facial dynamics transfer loss to improve their disentanglement. Let $\mathbf{I}_{i}$ and $\mathbf{I}_{j}$ be two frames randomly sampled from the same video. We extract their latent variables using the encoders, and transfer $\mathbf{I}_{i}$’s head pose onto $\mathbf{I}_{j}$ as $\hat{\mathbf{I}}_{j,\mathbf{z}^{pose}_{i}}=\mathcal{D}(\mathbf{V}^{app}_{j},\mathbf{z}^{id}_{j},\mathbf{z}^{pose}_{i},\mathbf{z}^{dyn}_{j})$ and $\mathbf{I}_{j}$’s facial motion onto $\mathbf{I}_{i}$ as $\hat{\mathbf{I}}_{i,\mathbf{z}^{dyn}_{j}}=\mathcal{D}(\mathbf{V}^{app}_{i},\mathbf{z}^{id}_{i},\mathbf{z}^{pose}_{i},\mathbf{z}^{dyn}_{j})$. The discrepancy loss $l_{consist}$ between $\hat{\mathbf{I}}_{j,\mathbf{z}^{pose}_{i}}$ and $\hat{\mathbf{I}}_{i,\mathbf{z}^{dyn}_{j}}$ is subsequently minimized. To reinforce the disentanglement between identity and motions, we add a face identity similarity loss $l_{cross\_id}$ for the cross-identity pose and facial motion transfer results. Let $\mathbf{I}_{s}$ and $\mathbf{I}_{d}$ be the video frames of two different subjects, we can transfer the motions of $\mathbf{I}_{d}$ onto $\mathbf{I}_{s}$ and obtain $\hat{\mathbf{I}}_{s,\mathbf{z}^{pose}_{d},\mathbf{z}^{dyn}_{d}}=\mathcal{D}(\mathbf{V}^{app}_{s},\mathbf{z}^{id}_{s},\mathbf{z}^{pose}_{d},\mathbf{z}^{dyn}_{d})$. Then, a cosine similarity loss between the deep face identity features [16] extracted from $\mathbf{I}_{s}$ and $\hat{\mathbf{I}}_{s,\mathbf{z}^{pose}_{d},\mathbf{z}^{dyn}_{d}}$ is applied. As we’ll show in the experiments, our new loss function deigns are crucial to achieve an effective factor disentanglement and facilitate the high-quality, lifelike talking face generation.
@@ 3.2 Holistic Facial Dynamics Generation with Diffusion Transformer
In order to achieve a seamless transition between adjacent windows, we incorporate the last $K$ frames of the audio feature and generated motions from the previous window as the condition of the current one. To summarize, our input condition can be denoted as $\mathbf{C}=[\mathbf{X}^{pre},\mathbf{A}^{pre};\mathbf{A},\mathbf{g},d,\mathbf{e}]$. All conditions are concatenated with noise along the temporal dimension as the input to the transformer.
@@ 4 Experiments
For face latent space learning, we use the public VoxCeleb2 dataset from [14] which contains talking face videos from about 6K subjects. We reprocess the dataset and discard the clips with multiple individuals and those of low quality using the method of [50]. For motion latent generation, we use an 8-layer transformer encoder with an embedding dim $512$ and head number $8$ as our diffusion network. The model is trained on VoxCeleb2 [14] and another high-resolution talk video dataset collected by us, which contains about 3.5K subjects. In our default setup, the model uses a forward-facing main gaze condition, an average head distance of all training videos, and an empty emotion offset condition. The CFG parameters are set to $\lambda_{\mathbf{A}}=0.5$ and $\lambda_{\mathbf{g}}=1.0$, and $50$ sampling steps are used. Our face latent model takes around 7 days of training on a 4 NVIDIA RTX A6000 GPUs workstation, and the diffusion transformer takes around 3 days. The total data used for training comprises approximately 500K clips, each lasting between 2 to 10 seconds. The parameter counts of our 3D-aided face latent model and diffusion transformer model are about 200M and 29M respectively.
@@
Inference speed.
Our method generates video frames of 512$\times$512 size at 45fps in the offline batch processing mode, and can support up to 40fps in the online streaming mode with a preceding latency of only 170ms , evaluated on a desktop PC with a single NVIDIA RTX 4090 GPU.
@@ 4.1 Quantitative Evaluation
  • Pose variation intensity. We further define a pose variation intensity score $\Delta P$ which is the average of the pose angle differences between adjacent frames. Averaged over all generated frames, $\Delta P$ provides an indication of the overall head motion intensity generated by a method.
  • Video quality. Following previous video generation works [69, 46], we use the Fréchet Video Distance (FVD) [57] to evaluate the generated video quality. We compute the FVD metric using sequences of 25 consecutive frames, at resolution of 224$\times$224.
@@
We compare our method with there existing audio-driven talking face generation methods: MakeItTalk [77], Audio2Head [62], and SadTalker [74].
Table 1: Quantitative comparison with previous methods on two benchmarks.
VoxCeleb2 OneMin-32
[…]
MakeItTalk 4.176 15.513 -0.051 0.210 -0.123 14.340 0.002 0.190 304.83
Audio2Head 6.172 8.470 0.246 0.260 5.992 8.211 0.205 0.239 209.77
SadTalker 5.843 8.813 0.441 0.275 5.501 8.850 0.383 0.252 214.51
Ours $\mathbb{8.841}$ $\mathbb{6.312}$ $\mathbb{0.468}$ $\mathbb{0.304}$ $\mathbb{7.957}$ $\mathbb{6.635}$ $\mathbb{0.465}$ $\mathbb{0.316}$ $\mathbb{105.88}$
Ours (10% data) 8.818 6.298 0.457 0.229 7.990 6.645 0.441 0.229 147.401
Real video 7.640 7.189 0.588 0.505 7.192 7.254 0.559 0.405 29.25
@@
For each audio input, we generate a single video for deterministic approaches, i.e., MakeItTalk and Audio2Head. For SadTalker and our method, we sample three videos for each audio and average the computed metrics. Since different pose representations are used by these methods, we re-extract the head poses from the generated frames to compute the pose-related metrics (i.e., CAPP and $\Delta P$). For the FVD metric, we use 2K 25-frame video clips of both the real videos and generated ones. For reference purpose, we also report the evaluated metrics of real videos.
Table 1 presents the results on the VoxCeleb2 and OneMin-32 benchmarks. Note that we did not evaluate the FVD on VoxCeleb2 as its video quality is varied and often low. On both benchmarks, our method achieves the best results among all methods on all evaluated metrics. In terms of audio-lip synchronization scores ($S_{C}$ and $S_{D}$), our method outperforms all others by a wide margin. Note that our method yields better scores than real videos, which is due to effect of the audio CFG (see Sec. 4.3). Our generated poses are better aligned with the audios especially on the OneMin-32 benchmark, as reflected by the CAPP scores. The head movements also exhibit the highest intensity according to $\Delta P$, although there’s still a gap to the intensity of real videos. Our FVD score is significantly lower than others, demonstrating the much higher video quality and realism of our results.
4.2 Qualitative Evaluation
Visual results.
Figure 1 presents some representative audio-driven talking face generation results of our method. Visually inspected, our method can generate high-quality video frames with vivid facial emotions. Moreover, it can generate human-like conversational behaviors, including sporadic shifts in eye gaze during speech and contemplation, as well as the natural and variable rhythm of eye blinking, among other nuances. We highly recommend that readers view our video results in the supplementary material to fully perceive the capabilities and output quality of our method.
Generation controllability.
Figure 3 shows our generated results under different control signals including main eye gaze, head distance, and emotion offset. Our model can well interpret these signals and produce talking face results that closely adhere to these specified parameters.
Disentanglement of face latents.
Figure A.1 shows that when applying the same motion latent sequences onto different subjects, our method effectively maintains both the distinct facial movements and the unique facial identities. This indicates the efficacy of our method in disentangling identity and motion. Figure A.2 further illustrates the effective disentanglement between head pose and facial dynamics. By holding one aspect constant and changing the other, the resulting images faithfully reflect the intended head and facial motions without interference.
Out-of-distribution generation.
Our method exhibits the capability to handle photo and audio inputs that fall outside the training distribution, such as artistic photos, singing audio clips, and non-English speech, as illustrated in Figure A.3.
Comparison with other methods.
Some visual examples from different methods are presented in Figure A.4 A.5 A.6 A.7. Our method outperforms the others in terms of the precise audio-lip synchronization and delivers much more vivid and natural facial dynamics and head movements.